# Efficient Transformed Gaussian Process State-Space Models for Non-Stationary High-Dimensional Dynamical Systems

### v3

@@ III-A Efficient Modeling for High-Dimensional State Space
Figure 2: (2(a)) Two warped GPs, see Eq. (10). $f_{1}(\cdot)$ effectively fits the data, while $f_{2}(\cdot)$ falls short. (2(b)) Independent GPs ($\tilde{f}_{1}(\cdot)$, $\tilde{f}_{2}(\cdot)$) vs our proposed ETGP (${f}_{1}(\cdot)$, ${f}_{2}(\cdot)$) depicted in Fig. 3. Both methods capture the underlying function reasonably well, though the GP $\tilde{f}_{1}(\cdot)$ exhibits slightly worse fit and overestimated uncertainty.
In Section II-C, we discussed the computational and modeling challenges that arise from using separate GPs for each output dimension of the transition function in high-dimensional state spaces. To address these limitations, [38] proposed warping a single shared GP using normalizing flows [44] to model each output dimension of the transition function:
@@
If ${\mathbb{G}}_{{{\bm{\theta}}}_{d}}(\cdot)$ is a Sinh-Arcsinh-Linear (SAL) flow [45, 46], we have:
@@
More complex flows can be achieved by stacking additional layers of SAL flow and/or incorporating more advanced deep neural network-based flows, such as RealNVP [47].
@@
However, the warped GP in Eq. (10) remains stationary; Also, being derived from a single shared GP, it may lack the expressiveness needed to capture complex, time-varying latent dynamics—resulting in suboptimal performance in practice [38]. An illustrative example is shown in Fig. 2(a), where the learned element-wise flow $\mathbb{G}_{\bm{\theta}_{1}}$ successfully warps $\tilde{f}$ to fit a simple target function $f_{1}$. In contrast, the warped function $f_{2}$ struggles to capture the time-varying and intricate structure, resulting in high uncertainty. For a more comprehensive comparison, Fig. 2(b) shows the results from both two independent GPs and the ETGP model introduced later in this work. While both capture the underlying structure, the stationary GP $\tilde{f}_{1}(\cdot)$ yields inferior fits and inflated uncertainty estimates.
To overcome these limitations in expressiveness, we draw inspiration from recent advances in adaptive parameterization [45, 48] and introduce non-stationarity and input dependence into the warped GPs. Specifically, we generate non-stationary processes $\{f_{d}(\cdot)\}_{d=1}^{d_{x}}$ by conditioning the flow parameters on the input through a shallow neural network $\textsc{nn}:\mathbb{R}^{d_{x}}\mapsto\Theta$ parameterized by weights $\mathbf{w}$. In the SSM context, this is formulated as [^3] Note that ${\mathbf{x}}_{t}$ can always be extended to accommodate control systems incorporating a deterministic control input $\bm{c}_{t}\in\mathbb{R}^{d_{c}}$ by augmenting the latent state to $[{\mathbf{x}}_{t}^{\top},\bm{c}_{t}^{\top}]^{\top}\in\mathbb{R}^{d_{x}+d_{c}}$. For brevity, however, we omit explicit reference to $\bm{c}_{t}$ in our notation throughout this paper.:
@@
where ${{\bm{\theta}}}_{t}=\{{{\bm{\theta}}}_{t,d}\}_{d=1}^{d_{x}}$. This formulation allows the flow parameters to vary with input at each time step $t$, enabling the model to capture local variations and time-varying dynamics. Additionally, note that the parameter amortization learning shown in Eq. (11) prevents the number of flow parameters from growing with time step $t$ while also improving parameter efficiency through weight sharing across dimensions.
We may further impose a prior $p_{\bm{\psi}}(\mathbf{w})$ over $\mathbf{w}$ with hyperparameter $\bm{\psi}$ (e.g., $\mathbf{w}\!\sim\!{\cal N}(\mathbf{w}|\bm{0},\operatorname{diag}(\bm{\psi}))$) to regularize the neural network’s parameters and mitigate overfitting issues. In summary, our modeling approach can be expressed mathematically as follows:
@@
${{\bm{\theta}}}_{t}=\textsc{nn}_{\mathbf{w}}({\mathbf{x}}_{t-1}),\ \ f_{d}(\cdot)={\mathbb{G}}_{{{\bm{\theta}}}_{t,d}}(\tilde{f}(\cdot)),\ d\!=\!1,...,{d_{x}},$ (12b)
where $\{f_{d}(\cdot)\}_{d=1}^{d_{x}}$ implicitly characterize $d_{x}$ random processes [43]. To better understand the modeling, Fig. 3 illustrates the corresponding graphical representation with $p_{\bm{\psi}}(\mathbf{w})=\delta(\mathbf{w}-\bar{\mathbf{w}})$, where $\delta(\cdot)$ is the Dirac delta measure and $\bar{\mathbf{w}}$ is a deterministic vector, and the following example extends the Example 1, providing a more concrete illustration.
@@
where ${{\bm{\theta}}}_{t}\!=\![\bm{\alpha}_{t}^{\top},\bm{\beta}_{t}^{\top}]^{\top}$, and $\bm{\alpha}_{t}\!=\!\{\alpha_{t,d}\}_{d=1}^{d_{x}}$, $\bm{\beta}_{t}\!=\!\{\beta_{t,d}\}_{d=1}^{d_{x}}$. An illustrative example of function fitting using the ETGP with linear flows is depicted in Fig. 2(b).
Figure 3: ETGP transition function in SSMs
It is noteworthy that this modeling strategy, referred to as the efficient transformed Gaussian process (ETGP) in the GP literature, was originally proposed to tackle large-scale multi-class classification problems [42]. Its efficiency arises from employing a single GP instead of one per class, ensuring that the computational cost of GP operations remains independent of the number of classes. This results in a substantial reduction in overall computation. In this work, we adapt the ETGP framework to model high-dimensional transition functions in SSMs. Notably, ETGP implicitly models $d_{x}$ dependent random processes [43], introducing output correlations that are often neglected in existing GPSSM formulations. We elaborate further on this below.
@@
If $p_{\bm{\psi}}(\mathbf{w})=\delta(\mathbf{w}-\bar{\mathbf{w}})$, where $\bar{\mathbf{w}}$ is a deterministic vector, then the ETGP defined in Example 3 becomes a $d_{x}$-dimensional dependent non-stationary Gaussian process with time-dependent covariance. Specifically, at arbitrary time steps $t$ and $t^{\prime}$, we have
@@
This result follows from the Gaussianity of the underlying process and the linearity of covariance, as detailed in Appendix C. ∎
@@
where the ETGP prior is defined by the base GP $\tilde{f}(\cdot)$, neural network weight prior $p_{\bm{\psi}}(\mathbf{w})$, and the normalizing flows, see Eq. (12). Unless stated otherwise, in this work we focus on the linear flow described in Example 3 for simplicity and illustration. This variant requires learning only two parameters per dimension, offering a more efficient alternative to more expressive but costly flow parameterizations. For simplicity, we also assume that the initial state prior $p({\mathbf{x}}_{0})$ is Gaussian and known.
@@
In contrast to existing GPSSMs, the $\{f_{d}(\cdot)\}_{d=1}^{d_{x}}$ in the ETGPSSM are dependent, as they are both transformed random processes derived from a shared GP and common weight priors (see Corollary 1). Furthermore, as these induced random processes are non-stationary (see Corollary 1), they are well-suited for modeling time-varying transition functions.
The primary objective in ETGPSSM, as in GPSSM, is the joint estimation of model hyperparameters $\bm{\vartheta}\!=\!\{{{\bm{\theta}}}_{gp},\bm{\psi},\mathbf{Q},\mathbf{R}\}$ and inference of both the ETGP and latent states. Given a sequence of observations ${\mathbf{y}}_{1:T}=\{{\mathbf{y}}_{t}\}_{t=1}^{T}$, with $T\in\mathbb{Z}_{+}$, the task is to estimate $\bm{\vartheta}$ and compute the posterior distributions $p(\bm{f}|{\mathbf{y}}_{1:T})$ for the ETGP and $p({\mathbf{x}}_{0:T}|{\mathbf{y}}_{1:T})$ for the latent states, where ${\mathbf{x}}_{0:T}=\{{\mathbf{x}}_{t}\}_{t=0}^{T}$.
However, the marginal likelihood $p(\mathbf{y}_{1:T}|\bm{\vartheta})$ is generally intractable, making both model evidence maximization and posterior inference computationally challenging. Furthermore, unlike standard GPs, which admit closed-form expressions for their prior and posterior (see Eqs. (2) and (3)), the ETGP defines an implicit stochastic process. As a result, both the marginal prior [^4] Obtaining $p(\bm{f})$ by marginalizing over the GP and $\mathbf{w}$ is generally intractable. and posterior distributions are analytically intractable [43], presenting significant challenges for inference. In the following subsections, we introduce a variational inference framework that enables joint learning of model hyperparameters and approximate Bayesian inference over the ETGP and latent states.
@@ III-B Approximate Bayesian Inference
To address the intractability of the marginal distribution $p({\mathbf{y}}_{1:T}|\bm{\vartheta})$, we employ variational inference. This approach maximizes a tractable evidence lower bound (ELBO) on the log-marginal likelihood, which is equivalent to minimizing the Kullback–Leibler (KL) divergence between the variational approximation and the true posterior [49]. Constructing the ELBO requires specifying both the model joint distribution and the variational distribution.
Following the existing GPSSM framework and building upon Eq. (16), we factorize the joint distribution as:
@@
where $p({\mathbf{x}}_{0})={\cal N}({\mathbf{x}}_{0}|\bm{0},\bm{I})$ and $p(\bm{f})$ denotes the ETGP prior (under slight abuse of notation). A common variational family in the GPSSM literature takes the form:
@@
While general, this formulation requires explicit parameterization of the variational distribution $q({\mathbf{x}}_{1:T}|\bm{f},{\mathbf{x}}_{0})$, which can introduce a large number of parameters and increase optimization complexity [37, 31].
@@
Compared to $\mathcal{L}_{2}$, the alternative bound $\mathcal{L}$ avoids direct parameterization of $q({\mathbf{x}}_{1:T}\mid\bm{f},{\mathbf{x}}_{0})$, helping reduce optimization overhead and yielding a tighter variational lower bound (see Proposition 1).
@@
Proof can be found in Appendix A
Theoretically, this tighter lower bound can improve inference accuracy and convergence [49]. We now proceed to derive the ELBO $\mathcal{L}$. As noted earlier, the lack of closed-form expressions for the ETGP’s prior and posterior prevents direct computation of $\mathcal{L}$ in function space. To overcome this, we leverage the generative construction of the ETGP—illustrated in Fig. 3—and infer the posterior distributions over $\tilde{f}$ and $\mathbf{w}$, which collectively characterize the ETGP posterior. Specifically, by explicitly representing $\tilde{f}$ and $\mathbf{w}$, the marginalized joint distribution becomes:
$p({\mathbf{y}}_{1:T},{\mathbf{x}}_{0},\mathbf{w},\tilde{f})=p({\mathbf{y}}_{1:T}|{\mathbf{x}}_{0},\mathbf{w},\tilde{f})p({\mathbf{x}}_{0})p_{\bm{\psi}}(\mathbf{w})p(\tilde{f}),$ (23)
where $p(\tilde{f})$ represents the GP prior and $p_{\bm{\psi}}(\mathbf{w})={\cal N}(\mathbf{w}|\bm{0},\operatorname{diag}(\bm{\psi}))$ is a Gaussian prior over the neural network weights, assuming a Bayesian neural network with diagonal covariance. Similarly, instead of using Eq. (21), we employ the following variational distribution:
@@
For the variational distribution $q(\tilde{f})$, we adopt a sparse GP approximation [40, 41], as defined in Eq. (8), which helps to reduce the computational complexity from $\mathcal{O}(T^{3})$ to $\mathcal{O}(M^{3})$.
@@
where the first terms (log-likelihood) can be analytically evaluated using the EnKF (detailed in Section III-C). The three KL divergence terms can also be computed in closed form, due to the Gaussian nature of the prior and variational distributions [49].
@@
The proof can be found in Appendix B. ∎
@@
The objective in Eq. (26) can also be derived from $\mathcal{L}_{2}$ under two specific assumptions (see Supplement A), akin to [37]. However, our derivation in this work is generic for arbitrary random processes and is assumption-free, offering greater simplicity, conceptual clarity, and broader applicability. For further details on the relationship between $\mathcal{L}$ and $\mathcal{L}_{2}$, we refer to Supplement  A.

### v4

@@ III-A Efficient Modeling for High-Dimensional State Space
Figure 2: Two-dimensional regression example comparing GP models. The non-stationary ETGP accurately captures local variations, outperforming stationary warped GPs and independent GPs.
In Section II-C, we discussed the computational and modeling challenges that arise from using separate GPs for each output dimension of the transition function in high-dimensional state spaces. To address these limitations, [lin2023towards_efficient] proposed warping a single shared GP using normalizing flows [kobyzev2020normalizing] to model each output dimension of the transition function:
@@
If ${\mathbb{G}}_{{{\bm{\theta}}}_{d}}(\cdot)$ is a Sinh-Arcsinh-Linear (SAL) flow [maronas2021transforming, rios2019compositionally], we have:
@@
More complex flows can be achieved by stacking additional layers of SAL flow and/or incorporating more advanced deep neural network-based flows, such as RealNVP [dinh2017density].
@@
However, the warped GP in Eq. (10) remains stationary; Also, being derived from a single shared GP, it may lack the expressiveness needed to capture complex latent dynamics—resulting in suboptimal performance in practice [lin2023towards_efficient]. A demonstrative example is provided as follows.
Example 3.
Consider a two-dimensional regression problem where the outputs are generated as follows:
$y_{1}$ $=3.5\cdot h(x),$
$y_{2}$ $=-1.5\cdot h(x)+5\sin(\pi x)+2\cos(2\pi x),$
where $h(x)=0.8+(x+0.2)\left(1-\frac{5}{1+\exp(-2x)}\right)+\varepsilon$ is a noisy kink function, and noise $\varepsilon\sim\mathcal{N}(0,0.25^{2})$. The function $y_{1}$ is linearly dependent on the noisy kink function, and $y_{2}$ introduces additional nonlinear, periodic components and a negative correlation with $y_{1}$. A dataset of 100 input-output pairs was generated with inputs in [4,2][-4,2], as shown in Fig. 2. The warped GPs, which are constructed by a base GP with an SE kernel and two SAL flows, were trained by maximizing the variational lower bound, see e.g. [maronas2021transforming, Section 4.1] for more details.
The results in Fig. 2(a) show that the learned element-wise SAL flow $\mathbb{G}_{\bm{\theta}_{1}}$ successfully warps the base GP $\tilde{f}$ into $f_{1}$, which fits $y_{1}$ well. In contrast, applying $\mathbb{G}_{\bm{\theta}_{2}}$ to the same base process yields $f_{2}$, which fails to capture the complex structure of $y_{2}$ due to the limited modeling flexibility of the input-independent SAL flows (see Example 2), resulting in high predictive uncertainty.
To overcome the limitations in expressiveness, we draw inspiration from recent advances in adaptive parameterization [maronas2021transforming, gu2023mamba] and introduce non-stationarity and input dependence into the warped GPs. Specifically, we generate non-stationary processes $\{f_{d}(\cdot)\}_{d=1}^{d_{x}}$ by conditioning the flow parameters on the input through a shallow neural network $\textsc{nn}:\mathbb{R}^{d_{x}}\mapsto\Theta$ parameterized by weights $\mathbf{w}$. In the SSM context, this is formulated as [^2] Note that ${\mathbf{x}}_{t}$ can always be extended to accommodate control systems incorporating a deterministic control input $\bm{c}_{t}\in\mathbb{R}^{d_{c}}$ by augmenting the latent state to ${\mathbf{x}}_{t}^{\top},\bm{c}_{t}^{\top}]^{\top}\in\mathbb{R}^{d_{x}+d_{c}}$. For brevity, however, we omit explicit reference to $\bm{c}_{t}$ in our notation throughout this paper.:
@@
where ${{\bm{\theta}}}_{t}=\{\theta_{t,d}\}_{d=1}^{d_{x}}$. This formulation allows the flow parameters to vary with input at each time step $t$, enabling the model to capture local variations and complex dynamics. Additionally, note that the parameter amortization learning shown in Eq. (11) prevents the number of flow parameters from growing with time step $t$ while also improving parameter efficiency through weight sharing across dimensions.
Figure 3: Non-stationary ETGP transition function in SSMs
We may further impose a prior $p_{\bm{\psi}}(\mathbf{w})$ over $\mathbf{w}$ with hyperparameter $\bm{\psi}$ (e.g., $\mathbf{w}\!\sim\!{\cal N}(\mathbf{w}|\bm{0},\operatorname{diag}(\bm{\psi}))$) to regularize the neural network’s parameters and mitigate overfitting issues. In summary, this efficient transformed GP modeling can be mathematically expressed as follows:
@@
${{\bm{\theta}}}_{t}=\textsc{nn}_{\mathbf{w}}({\mathbf{x}}_{t-1}),\ \ f_{d}(\cdot)={\mathbb{G}}_{\theta_{t,d}}(\tilde{f}(\cdot)),\ d\!=\!1,...,{d_{x}},$ (12b)
where $\{f_{d}(\cdot)\}_{d=1}^{d_{x}}$ implicitly characterize $d_{x}$ random processes [ma2019variational]. To better understand the modeling, Fig. 3 illustrates the corresponding graphical representation with $p_{\bm{\psi}}(\mathbf{w})=\delta(\mathbf{w}-\bar{\mathbf{w}})$, where $\delta(\cdot)$ is the Dirac delta measure and $\bar{\mathbf{w}}$ is a deterministic vector. The following example extends the Example 1, providing a more concrete illustration.
@@
where ${{\bm{\theta}}}_{t}\!=\![\bm{\alpha}_{t}^{\top},\bm{\beta}_{t}^{\top}]^{\top}$, and $\bm{\alpha}_{t}\!=\!\{\alpha_{t,d}\}_{d=1}^{d_{x}}$, $\bm{\beta}_{t}\!=\!\{\beta_{t,d}\}_{d=1}^{d_{x}}$.
An illustrative two-dimensional regression example using the resulting transformed GPs, ${f}_{1}$ and ${f}_{2}$, is shown in Fig. 2(b), where the neural network weights follow a Dirac delta measure. The training data consist of 100 input–output pairs consistent with Example 3. For comparison, results from two independent GPs, $\tilde{f}_{1}$ and $\tilde{f}_{2}$, with SE kernels are also presented. Both models were trained by maximizing their respective variational lower bounds [hensman2013gaussian, maronas2022efficient].
While both models capture the general structure of the target functions, the independent GP $\tilde{f}_{2}$ exhibits a poorer fit and overestimates uncertainty for the more complex $y_{2}$, primarily due to its stationary SE kernel’s inability to adapt to local variations. In contrast, the ETGP achieves a notably better fit for these complex patterns through its input-dependent linear flow, which induces non-stationary behavior. Moreover, the ETGP shares a single latent GP across outputs, offering improved parameter efficiency and scalability compared to modeling each output with a separate GP, as discussed in Section II-C.
It is noteworthy that this efficient transformed Gaussian process (ETGP) modeling strategy was originally proposed to tackle large-scale multi-class classification problems [maronas2022efficient]. Its efficiency arises from employing a single GP instead of one per class, ensuring that the computational cost of GP operations remains independent of the number of classes. This results in a substantial reduction in overall computation. In this work, we adapt this framework to model high-dimensional transition functions in SSMs. Notably, ETGP implicitly models $d_{x}$ dependent random processes [ma2019variational], introducing output correlations that are often neglected in existing GPSSM formulations. We elaborate further on this below.
@@
If $p_{\bm{\psi}}(\mathbf{w})=\delta(\mathbf{w}-\bar{\mathbf{w}})$, where $\bar{\mathbf{w}}$ is a deterministic vector, then the ETGP defined in Example 4 becomes a $d_{x}$-dimensional dependent non-stationary Gaussian process with an input-dependent covariance function. Specifically, at arbitrary time steps $t$ and $t^{\prime}$, we have
@@
This result follows from the Gaussianity of the underlying process and the linearity of covariance, as detailed in Appendix A. ∎
@@
[…]
@@
In contrast to existing GPSSMs with independent, stationary transitions, the ETGPSSM yields dependent and non-stationary transition functions $\{f_{d}(\cdot)\}_{d=1}^{d_{x}}$ due to their shared base GP and weight priors. In the case of linear input-dependent flows, the transition function remains a GP—like a GPSSM—but enriched with dependence and non-stationarity, see Corollary 1, offering greater modeling flexibility while retaining GP interpretability.
The primary objective in ETGPSSM, as in GPSSM, is the joint estimation of model hyperparameters $\bm{\vartheta}\!=\!\{{{\bm{\theta}}}_{gp},\bm{\psi},\mathbf{Q},\mathbf{R}\}$ and inference of both the ETGP and latent states. Given a sequence of observations ${\mathbf{y}}_{1:T}$, the task is to estimate $\bm{\vartheta}$ and compute the posterior distributions $p(\bm{f}|{\mathbf{y}}_{1:T})$ for the ETGP and $p({\mathbf{x}}_{0:T}|{\mathbf{y}}_{1:T})$ for the latent states, where ${\mathbf{x}}_{0:T}=\{{\mathbf{x}}_{t}\}_{t=0}^{T}$.
However, the marginal likelihood $p(\mathbf{y}_{1:T}|\bm{\vartheta})$ is generally intractable, making both model evidence maximization and posterior inference computationally challenging. Furthermore, unlike GPs, which admit closed-form expressions for their prior and (approximate) posterior (see Eqs. (2), (3) & (8)), the ETGP defines an implicit stochastic process. As a result, both the prior and posterior are analytically intractable [ma2019variational], presenting significant challenges for inference. In the following subsections, we introduce a variational inference framework that enables joint learning of model hyperparameters and approximate Bayesian inference over the ETGP and latent states.
@@ III-B Approximate Bayesian Inference
Variational inference proceeds by maximizing a more tractable evidence lower bound (ELBO) on the log-marginal likelihood, which is equivalent to minimizing the Kullback–Leibler (KL) divergence between the variational approximation and the true posterior [theodoridis2020machine]. Constructing the ELBO requires specifying both the model joint distribution and the variational distribution.
Following the existing GPSSM framework, we factorize the joint distribution of the ETGPSSM defined in Eq. (16) as:
@@
where $p({\mathbf{x}}_{0})={\cal N}({\mathbf{x}}_{0}|\bm{0},\bm{I})$, and under slight abuse of notation [^3] In general a random process does not have density function., $p(\bm{f})$ denotes the ETGP prior, which is generally intractable because it requires marginalizing over the base GP $\tilde{f}$ and the neural network weights $\mathbf{w}$. A common variational family in the GPSSM literature takes the form:
@@
While general, this formulation requires explicit parameterization of the variational distribution $q({\mathbf{x}}_{1:T}|\bm{f},{\mathbf{x}}_{0})$, which can introduce a large number of parameters and increase optimization complexity [lin2023ensemble, ialongo2019overcoming].
@@
Compared to $\mathcal{L}_{2}$, the alternative bound $\mathcal{L}$ avoids direct parameterization of $q({\mathbf{x}}_{1:T}|\bm{f},{\mathbf{x}}_{0})$, helping reduce optimization overhead and yielding a tighter variational lower bound (see Proposition 1).
@@
Proof can be found in Appendix B
Theoretically, this tighter lower bound can improve inference accuracy and convergence [theodoridis2020machine]. Next, we proceed to derive the ELBO $\mathcal{L}$.
As noted earlier, the lack of closed-form expressions for the ETGP’s prior and posterior prevents direct computation of $\mathcal{L}$ in function space. To overcome this, we leverage the generative construction of the ETGP—illustrated in Fig. 3—and infer the posterior distributions over $\tilde{f}$ and $\mathbf{w}$, which collectively characterize the ETGP posterior. Specifically, by explicitly representing $\tilde{f}$ and $\mathbf{w}$, the marginalized joint distribution becomes:
[…]
where p(f~)p($\tilde${f}) represents the GP prior and $p_{\bm{\psi}}(\mathbf{w})={\cal N}(\mathbf{w}|\bm{0},\operatorname{diag}(\bm{\psi}))$ is a Gaussian prior over the neural network weights, assuming a Bayesian neural network with diagonal covariance. Similarly, instead of using Eq. (21), we employ the following variational distribution:
@@
For the variational distribution q(f~)q($\tilde${f}), we adopt a sparse GP approximation [titsias2009variational, hensman2013gaussian], as defined in Eq. (8), which helps to reduce the computational complexity from $\mathcal{O}(T^{3})$ to $\mathcal{O}(M^{3})$.
@@
where the first terms (log-likelihood) can be analytically evaluated using the EnKF (detailed in Section III-C). The three KL divergence terms can also be computed in closed form, due to the Gaussian nature of the prior and variational distributions [theodoridis2020machine].
@@
The proof can be found in Appendix C. ∎
@@
The objective in Eq. (26) can also be derived from $\mathcal{L}_{2}$ under two specific assumptions (see Supplement A), akin to [lin2023ensemble]. However, our derivation in this work is generic for arbitrary random processes and is assumption-free, offering greater simplicity, conceptual clarity, and broader applicability. For further details on the relationship between $\mathcal{L}$ and $\mathcal{L}_{2}$, we refer the reader to Supplement A.